You have $30 to spend for a sweatshirt. the sales tax is 6% What is the most the sweatshirt can cost? Write an equation and solve.
Answer by HyperBrain(694)

Let x=the cost of the shirt
Then,
0.06x=the tax of the shirt
Thus, if the total cost---the shirt and the tax---have to reach the maximum amount, $30.00,
x+0.06x=30
100*(x+0.06x)=100*(30)
100x+6x=3000
106x=3000
x=$28.30

therefore, he could only buy a shirt with amaximum price of $28.30!